# Pick up the Pace: A Parameter-Free Optimizer for Lifelong Reinforcement Learning

### v1

@@ Abstract
A key challenge in lifelong reinforcement learning (RL) is the loss of plasticity, where previous learning progress hinders an agent’s adaptation to new tasks. While regularization and resetting can help, they require precise hyperparameter selection at the outset and environment-dependent adjustments. Building on the principled theory of online convex optimization, we present a parameter-free optimizer for lifelong RL, called Pace, which requires no tuning or prior knowledge about the distribution shifts. Extensive experiments on Procgen, Atari, and Gym Control environments show that Pace works surprisingly well—mitigating loss of plasticity and rapidly adapting to challenging distribution shifts—despite the underlying optimization problem being nonconvex and nonstationary.
@@ 1 Introduction
From the optimization perspective, the above issues might be attributed to the lack of stability under gradient descent. That is, the weights of the agent’s parameterized policy can drift far away from the origin (or a good initialization), leading to a variety of undesirable behaviors. [^1] Such as the inactivation of many neurons, due to the ReLU activation function (Abbas et al., 2023; Sokar et al., 2023). Fitting this narrative, it has been shown that simply adding a $L_{2}$ regularizer to the optimization objective (Kumar et al., 2023) or periodically resetting the weights (Dohare et al., 2021; Asadi et al., 2023; Sokar et al., 2023; Ahn et al., 2024) can help mitigate the problem. However, a particularly important limitation is their use of hyperparameters, such as the magnitude of the regularizer and the resetting frequency. Good performance hinges on the suitable environment-dependent hyperparameter, but how can one confidently choose that before interacting with the environment? The classical cross-validation approach would violate the one-shot nature of lifelong RL (and online learning in general; see Chapter 1 of Orabona, 2023), since it is impossible to experience the same environment multiple times. This leads to the contributions of the present work.
@@
  • Algorithm: Pace Building on a series of results in OCO (Cutkosky & Orabona, 2018; Cutkosky, 2019; Cutkosky et al., 2023; Zhang et al., 2024b), we propose a (hyper)-parameter-free optimizer for lifelong RL, called Pace (Parameter-free Adaption for Continual Environments). Intuitively, the idea is a refinement of regularization: instead of manually selecting the magnitude of regularization beforehand, Pace chooses that in an online, data-dependent manner. From the perspective of OCO theory, Pace is insensitive to its own hyperparameter, which means that no hyperparameter tuning is necessary in practice. Furthermore, as an optimization approach to lifelong RL, Pace is compatible with any policy parameterization.
  • Experiment Using Proximal Policy Optimization (PPO), we conduct comprehensive experiments on the instantiation of Pace called Pace PPO. A diverse range of lifelong RL environments are tested (based on Procgen, Atari, and Gym Control), with considerably larger scale than prior works. In settings where existing approaches (Abbas et al., 2023; Kumar et al., 2023) struggle, we find that Pace PPO
@@
    Such findings might be surprising: the theoretical advantage of Pace is motivated by the convexity in OCO, but lifelong RL is both nonconvex and nonstationary in terms of optimization.
@@
Section 2 surveys the basics of lifelong RL. Section 3 introduces our parameter-free algorithm Pace, and experiments are presented in Section 4. We defer the discussion of related works and results to Section 5. Finally, Section 6 concludes the paper.
@@ 3 Method
Figure 2: Visualization of Pace’s key idea.
Inspired by (Cutkosky et al., 2023), we study lifelong RL by exploiting its connection to Online Convex Optimization (OCO; Zinkevich, 2003). The latter is a classical theoretical problem in online learning, and much effort has been devoted to designing parameter-free algorithms that require minimum tuning or prior knowledge (Streeter & Mcmahan, 2012; McMahan & Orabona, 2014; Orabona & Pál, 2016; Foster et al., 2017; Cutkosky & Orabona, 2018; Mhammedi & Koolen, 2020; Chen et al., 2021; Jacobsen & Cutkosky, 2022). The surprising observation of Cutkosky et al. (2023) is that several algorithmic ideas closely tied to the convexity of OCO can actually improve the nonconvex deep learning training, suggesting certain notions of “near convexity” on its loss landscape. We find that lifelong RL (which is both nonconvex and nonstationary in terms of optimization) exhibits a similar behavior, therefore a particularly strong algorithm (named Pace) can be obtained from principled results in parameter-free OCO. Let us start from the background.
@@
Pace for lifelong RL
Now back to lifelong RL. As we discussed, a fundamental challenge here is the excessive drifting of the weights $\theta_{t}$, and this can be fixed by enforcing the proximity to a good reference point $\theta_{\mathrm{ref}}$. Different from existing approaches like $L_{2}$ regularization and resetting, parameter-free OCO provides a principled solution to this problem without hyperparameter-tuning. Naming this algorithm as Pace, we present its generic template as Algorithm 1, which calls Algorithm 2 (a one-dimensional scale tuner) as the key subroutine.
From the technical perspective, Pace assembles three techniques in parameter-free OCO: the direction-magnitude decomposition from (Cutkosky & Orabona, 2018), the additive aggregation from (Cutkosky, 2019), and the $\mathrm{erfi}$ potential function from (Zhang et al., 2024b). In addition, we follow (Cutkosky et al., 2023) for the recommended default parameter setup.
Algorithm 1 Pace: Parameter-free Adaption for Continual Environments.
@@
Algorithm 2 1D Discounted Tuner of Pace.
@@
  • First, Pace is a meta-algorithm that operates on top of a “default” optimizer Base. It can simply be gradient descent with a constant learning rate, or Adam (Kingma & Ba, 2014) as in our experiments. Applying Base alone would be equivalent to enforcing the scaling parameter $S_{t+1}\equiv 1$ in Pace, but this would suffer from the drifting of $\theta_{t+1}^{\mathrm{Base}}$ (and thus, the weight $\theta_{t+1}$).
  • To fix this issue, Pace uses the tuner (Algorithm 2) to select the scaling parameter $S_{t+1}$, making it data-dependent. Typically $S_{t+1}$ is within $[0,1]$ (see Figure 12 to 14), therefore essentially, we define the updated weight $\theta_{t+1}$ as a convex combination of the Base’s weight $\theta_{t}^{\mathrm{Base}}$ and the reference point $\theta_{\mathrm{ref}}$,
@@
Despite its nested structure, Pace can actually be seen as a parameter-free refinement of $L_{2}$ regularization (Kumar et al., 2023). To concretely explain this intuition, let us consider the following two optimization dynamics.
@@
  • Alternatively, consider applying the following simplification of Pace on the same policy gradient sequence $\{g_{t}\}$: ($i$) Base is still gradient descent with learning rate $\eta$; ($ii$) there is just one discount factor $\beta$; and ($iii$) the one-dimensional tuner (Algorithm 2) is replaced by the $\beta$-discounted gradient descent with learning rate $\alpha$, i.e., $S_{t+1}=\beta S_{t}-\alpha h_{t}$. In this case, we have
@@
The main message here is that under a simplified setting, Pace is almost equivalent to $L_{2}$ regularization. The latter requires choosing the hyperparameters $\lambda$ and $\eta$, and similarly, the above simplified Pace requires choosing $\beta$ and $\eta$. Going beyond this simplification, the actual Pace removes the tuning of $\beta$ using aggregation, and the tuning of $\eta$ using the $\mathrm{erfi}$ decision rule.
@@
Although Pace is called “parameter-free”, it still needs the $\beta$-grid, the constant $\varepsilon$ and the algorithm Base as inputs. The idea is that Pace is particularly insensitive to such choices, as supported by the OCO theory. As the result, the generic default values recommended by Cutkosky et al. (2023) are sufficient in practice. We note that those are proposed for training supervised deep learning models, thus should be agnostic to the lifelong RL applications we consider.
@@ 4 Experiment
Table 1: Cumulative sum of mean episode reward for Pace PPO, Adam PPO, and CReLU on Procgen, Atari, and Gym Control environments. Rewards are scaled by $10^{5}$; higher is better.
Environment Adam PPO CReLU Pace PPO (Ours)
@@
Does Pace experience the common pitfalls of loss of plasticity? Does it rapidly adapt to distribution shifts? To answer these questions, we test Pace in empirical RL benchmarks such as vision-based games and physics-based control environments in lifelong settings (Figure 3). Specifically, we instantiate PPO with two different optimizers: Adam with constant learning rate for baseline comparison, and Pace for our proposed method (with exactly the same Adam as the input Base). We also test Adam PPO with concatenated ReLU activations (CReLU; Shang et al., 2016), previously shown to mitigate loss of plasticity in certain deep RL settings (Abbas et al., 2023). Our numerical results are summarized in Table 1. Across every lifelong RL setting, we observe substantial improvements in the cumulative episode reward by using Pace PPO compared to Adam PPO or CReLU. Below are the details, with more in Appendix E.
@@
We first evaluate on OpenAI Procgen, a suite of 16 procedurally generated game environments (Cobbe et al., 2019). We introduce distribution shifts by sampling a new procedurally generated level of the current game every 2 million time steps, treating each level as a distinct task.
Figure 4: Reward in the lifelong Procgen environments for StarPilot, Dodgeball, Fruitbot, and Chaser. The plots demonstrate a steady loss of plasticity in agents using Adam PPO and CReLU, characterized by their inability to maintain performance through succesive Procgen levels. In contrast, Pace exhibits robustness against this loss of plasticity, quickly achieving high performance with each new task.
We evaluate game environments including StarPilot, Dodgeball, Fruitbot, and Chaser. In all of these environments, we observe in Figure 4 that both Adam PPO and CReLU encounter a continually degrading loss of plasticity as these distribution shifts are introduced. In contrast, Pace PPO avoids this loss of plasticity, which contributes to its rapid reward increase when adapting to new levels. In the cumulative reward across all the Procgen levels, Pace PPO reveals normalized average improvements of 3,212.42% and 120.88% over Adam PPO and CReLU respectively (see Table 1). For later levels, in all games, Pace PPO’s reward does not decline as sharply as the baselines, potentially indicating forward transfer of skills from one level to the next.
@@
The Arcade Learning Environment (ALE) Atari 2600 benchmark is a collection of classic arcade games designed to assess reinforcement learning agents’ performance across a range of diverse gaming scenarios (Bellemare et al., 2012). We introduce distribution shifts by switching to a new Atari game every 4 million timesteps, where each game switch introduces a new task. This benchmark is more challenging compared to OpenAI Procgen: it requires the agent to handle distribution shifts in both the input (state) and the target (reward).
Figure 5: Reward in the lifelong Atari environments, across games with action spaces of 6 and 9. These plots demonstrate that Pace PPO rapidly adapts to new tasks, in contrast to the Adam PPO and CReLU which struggle to achieve high reward, indicating mild loss of plasticity.
In this experiment, we assessed two online settings distinguished by games with action spaces of 6 and 9. From Figure 5, both Adam PPO and CReLU sometimes failed to learn in certain games. In contrast, Pace PPO shows a substantial increase in reward over different games compared to the baselines. For example, during the first 12 million steps (3 games) in Atari 6, Pace PPO not only achieves a significantly higher mean reward but also demonstrates rapid reward increase. Over both experiment settings, Pace PPO shows an average normalized improvement of 329.73% over Adam PPO and 68.71% over CReLU (Table 1). In rare instances, such as the last 2 million steps of Atari 6, CReLU performs comparably to Pace PPO. This observation aligns with findings from (Abbas et al., 2023), which noted that CReLU tends to avoid plasticity loss in continual Atari setups.
@@
We use the CartPole-v1 and Acrobot-v1 environments from the Gym Classic Control suite, along with LunarLander-v2 from Box2d Control. To introduce distribution shifts, a method outlined by Mendez et al. (2020) involves periodically altering the environment dynamics. Although such distribution shifts pose only mild challenges for robust methods like PPO with Adam (Appendix B). We instead implement a more challenging form of distribution shift. Every 200 steps we perturb each observation dimension with random noise within a range of ±$\pm 2$, treating each perturbation phase as a distinct task.
Figure 6: Reward performance across CartPole, Acrobot, and LunarLander Gym Control tasks. In these environments, both Adam PPO and CReLU experience extreme plasticity loss, failing to recover after the initial distribution shift. Conversely, Pace PPO successfully avoids such plasticity loss, rapidly adapting to each extreme distribution shift.
Here (Figure 6), we notice a peculiar behavior after introducing the first distribution shift in both Adam PPO and CReLU: policy collapse. We describe this as an extreme form of loss of plasticity. Surprisingly, Pace PPO remains resistant to these extreme distribution shifts. As we see in the Acrobot experiment, Pace PPO shows minimal to no policy damage after the first few distribution shifts, whereas Adam PPO and CReLU are unable to recover a policy at all. Across the three Gym Control environments, Pace PPO shows an average normalized improvement of 204.18% over Adam PPO and 1044.24% over CReLU (Table 1).

### v2

# Fast Trac: A Parameter-Free Optimizer for Lifelong Reinforcement Learning

@@ Abstract
A key challenge in lifelong reinforcement learning (RL) is the loss of plasticity, where previous learning progress hinders an agent’s adaptation to new tasks. While regularization and resetting can help, they require precise hyperparameter selection at the outset and environment-dependent adjustments. Building on the principled theory of online convex optimization, we present a parameter-free optimizer for lifelong RL, called Trac, which requires no tuning or prior knowledge about the distribution shifts. Extensive experiments on Procgen, Atari, and Gym Control environments show that Trac works surprisingly well—mitigating loss of plasticity and rapidly adapting to challenging distribution shifts—despite the underlying optimization problem being nonconvex and nonstationary.
@@ 1 Introduction
From the optimization perspective, the above issues might be attributed to the lack of stability under gradient descent. That is, the weights of the agent’s parameterized policy can drift far away from the origin (or a good initialization), leading to a variety of undesirable behaviors. [^1] Such as the inactivation of many neurons, due to the ReLU activation function (Abbas et al., 2023; Sokar et al., 2023). Fitting this narrative, it has been shown that simply adding a $L_{2}$ regularizer to the optimization objective (Kumar et al., 2023) or periodically resetting the weights (Dohare et al., 2021; Asadi et al., 2024; Sokar et al., 2023; Ahn et al., 2024) can help mitigate the problem. However, a particularly important limitation is their use of hyperparameters, such as the magnitude of the regularizer and the resetting frequency [^2] Indeed, hyperparameter selection, in general, is a well-known problem in lifelong as well as continual learning settings (De Lange et al., 2021).. Good performance hinges on the suitable environment-dependent hyperparameter, but how can one confidently choose that before interacting with the environment? The classical cross-validation approach would violate the one-shot nature of lifelong RL (and online learning in general; see Chapter 1 of Orabona, 2023), since it is impossible to experience the same environment multiple times. This leads to the contributions of the present work.
@@
  • Algorithm: Trac Building on a series of results in OCO (Cutkosky & Orabona, 2018; Cutkosky, 2019; Cutkosky et al., 2023; Zhang et al., 2024b), we propose a (hyper)-parameter-free optimizer for lifelong RL, called Trac (AdapTive RegularizAtion in Continual environments). Intuitively, the idea is a refinement of regularization: instead of manually selecting the magnitude of regularization beforehand, Trac chooses that in an online, data-dependent manner. From the perspective of OCO theory, Trac is insensitive to its own hyperparameter, which means that no hyperparameter tuning is necessary in practice. Furthermore, as an optimization approach to lifelong RL, Trac is compatible with any policy parameterization.
  • Experiment Using Proximal Policy Optimization (PPO), we conduct comprehensive experiments on the instantiation of Trac called Trac PPO. A diverse range of lifelong RL environments are tested (based on Procgen, Atari, and Gym Control), with considerably larger scale than prior works. In settings where existing approaches (Abbas et al., 2023; Kumar et al., 2023) struggle, we find that Trac PPO
@@
    Such findings might be surprising: the theoretical advantage of Trac is motivated by the convexity in OCO, but lifelong RL is both nonconvex and nonstationary in terms of optimization.
@@
Section 2 surveys the basics of lifelong RL. Section 3 introduces our parameter-free algorithm Trac, and experiments are presented in Section 4. We defer the discussion of related works and results to Section 5. Finally, Section 6 concludes the paper.
@@ 3 Method
Figure 2: Visualization of Trac’s key idea.
Inspired by (Cutkosky et al., 2023), we study lifelong RL by exploiting its connection to Online Convex Optimization (OCO; Zinkevich, 2003). The latter is a classical theoretical problem in online learning, and much effort has been devoted to designing parameter-free algorithms that require minimum tuning or prior knowledge (Streeter & Mcmahan, 2012; McMahan & Orabona, 2014; Orabona & Pál, 2016; Foster et al., 2017; Cutkosky & Orabona, 2018; Mhammedi & Koolen, 2020; Chen et al., 2021; Jacobsen & Cutkosky, 2022). The surprising observation of Cutkosky et al. (2023) is that several algorithmic ideas closely tied to the convexity of OCO can actually improve the nonconvex deep learning training, suggesting certain notions of “near convexity” on its loss landscape. We find that lifelong RL (which is both nonconvex and nonstationary in terms of optimization) exhibits a similar behavior, therefore a particularly strong algorithm (named Trac) can be obtained from principled results in parameter-free OCO. Let us start from the background.
@@
Trac for lifelong RL
Now back to lifelong RL. As we discussed, a fundamental challenge here is the excessive drifting of the weights $\theta_{t}$, and this can be fixed by enforcing the proximity to a good reference point $\theta_{\mathrm{ref}}$. Different from existing approaches like $L_{2}$ regularization and resetting, parameter-free OCO provides a principled solution to this problem without hyperparameter-tuning. Naming this algorithm as Trac, we present its generic template as Algorithm 1, which calls Algorithm 2 (a one-dimensional scale tuner) as the key subroutine.
From the technical perspective, Trac assembles three techniques in parameter-free OCO: the direction-magnitude decomposition from (Cutkosky & Orabona, 2018), the additive aggregation from (Cutkosky, 2019), and the $\mathrm{erfi}$ potential function from (Zhang et al., 2024b). In addition, we follow (Cutkosky et al., 2023) for the recommended default parameter setup.
Algorithm 1 Trac: Parameter-free Adaption for Continual Environments.
@@
Algorithm 2 1D Discounted Tuner of Trac.
@@
  • First, Trac is a meta-algorithm that operates on top of a “default” optimizer Base. It can simply be gradient descent with a constant learning rate, or Adam (Kingma & Ba, 2014) as in our experiments. Applying Base alone would be equivalent to enforcing the scaling parameter $S_{t+1}\equiv 1$ in Trac, but this would suffer from the drifting of $\theta_{t+1}^{\mathrm{Base}}$ (and thus, the weight $\theta_{t+1}$).
  • To fix this issue, Trac uses the tuner (Algorithm 2) to select the scaling parameter $S_{t+1}$, making it data-dependent. Typically $S_{t+1}$ is within $[0,1]$ (see Figure 13 to 15), therefore essentially, we define the updated weight $\theta_{t+1}$ as a convex combination of the Base’s weight $\theta_{t}^{\mathrm{Base}}$ and the reference point $\theta_{\mathrm{ref}}$,
@@
Despite its nested structure, Trac can actually be seen as a parameter-free refinement of $L_{2}$ regularization (Kumar et al., 2023). To concretely explain this intuition, let us consider the following two optimization dynamics.
@@
  • Alternatively, consider applying the following simplification of Trac on the same policy gradient sequence $\{g_{t}\}$: ($i$) Base is still gradient descent with learning rate $\eta$; ($ii$) there is just one discount factor $\beta$; and ($iii$) the one-dimensional tuner (Algorithm 2) is replaced by the $\beta$-discounted gradient descent with learning rate $\alpha$, i.e., $S_{t+1}=\beta S_{t}-\alpha h_{t}$. In this case, we have
@@
The main message here is that under a simplified setting, Trac is almost equivalent to $L_{2}$ regularization. The latter requires choosing the hyperparameters $\lambda$ and $\eta$, and similarly, the above simplified Trac requires choosing $\beta$ and $\eta$. Going beyond this simplification, the actual Trac removes the tuning of $\beta$ using aggregation, and the tuning of $\eta$ using the $\mathrm{erfi}$ decision rule.
@@
Although Trac is called “parameter-free”, it still needs the $\beta$-grid, the constant $\varepsilon$ and the algorithm Base as inputs. The idea is that Trac is particularly insensitive to such choices, as supported by the OCO theory. As the result, the generic default values recommended by Cutkosky et al. (2023) are sufficient in practice. We note that those are proposed for training supervised deep learning models, thus should be agnostic to the lifelong RL applications we consider.
@@ 4 Experiment
Table 1: Cumulative sum of mean episode reward for Trac PPO, Adam PPO, and CReLU on Procgen, Atari, and Gym Control environments. Rewards are scaled by $10^{5}$; higher is better.
Environment Adam PPO CReLU Trac PPO (Ours)
@@
Does Trac experience the common pitfalls of loss of plasticity? Does it rapidly adapt to distribution shifts? To answer these questions, we test Trac in empirical RL benchmarks such as vision-based games and physics-based control environments in lifelong settings (Figure 3). Specifically, we instantiate PPO with two different optimizers: Adam with constant learning rate for baseline comparison, and Trac for our proposed method (with exactly the same Adam as the input Base). We also test Adam PPO with concatenated ReLU activations (CReLU; Shang et al., 2016), previously shown to mitigate loss of plasticity in certain deep RL settings (Abbas et al., 2023). Our numerical results are summarized in Table 1. Across every lifelong RL setting, we observe substantial improvements in the cumulative episode reward by using Trac PPO compared to Adam PPO or CReLU. Below are the details, with more in the Appendix.
@@
We first evaluate on OpenAI Procgen, a suite of 16 procedurally generated game environments (Cobbe et al., 2020). We introduce distribution shifts by sampling a new procedurally generated level of the current game every 2 million time steps, treating each level as a distinct task.
Figure 4: Reward in the lifelong Procgen environments for StarPilot, Dodgeball, Fruitbot, and Chaser. There is a steady loss of plasticity in agents using Adam PPO and CReLU, characterized by their inability to maintain performance through succesive Procgen levels. In contrast, Trac avoids this loss of plasticity, quickly achieving high performance with each new task.
We evaluate game environments including StarPilot, Dodgeball, Fruitbot, and Chaser. In all of these environments, we observe in Figure 4 that both Adam PPO and CReLU encounter a continually degrading loss of plasticity as these distribution shifts are introduced. In contrast, Trac PPO avoids this loss of plasticity, which contributes to its rapid reward increase when adapting to new levels. In the cumulative reward across all the Procgen levels, Trac PPO reveals normalized average improvements of 3,212.42% and 120.88% over Adam PPO and CReLU respectively (see Table 1). For later levels, in all games, Trac PPO’s reward does not decline as sharply as the baselines, potentially indicating positive transfer of skills from one level to the next.
@@
The Arcade Learning Environment (ALE) Atari 2600 benchmark is a collection of classic arcade games designed to assess reinforcement learning agents’ performance across a range of diverse gaming scenarios (Bellemare et al., 2013). We introduce distribution shifts by switching to a new Atari game every 4 million timesteps, where each game switch introduces a new task. This benchmark is more challenging compared to OpenAI Procgen: it requires the agent to handle distribution shifts in both the input (state) and the target (reward).
Figure 5: Reward in the lifelong Atari environments, across games with action spaces of 6 and 9. Trac PPO rapidly adapts to new tasks, in contrast to the Adam PPO and CReLU which struggle to achieve high reward, indicating mild loss of plasticity.
In this experiment, we assessed two online settings distinguished by games with action spaces of 6 and 9. From Figure 5, both Adam PPO and CReLU sometimes failed to learn in certain games. In contrast, Trac PPO shows a substantial increase in reward over different games compared to the baselines. For example, during the first 12 million steps (3 games) in Atari 6, Trac PPO not only achieves a significantly higher mean reward but also demonstrates rapid reward increase. Over both experiment settings, Trac PPO shows an average normalized improvement of 329.73% over Adam PPO and 68.71% over CReLU (Table 1). In rare instances, such as the last 2 million steps of Atari 6, CReLU performs comparably to Trac PPO. This observation aligns with findings from (Abbas et al., 2023), which noted that CReLU tends to avoid plasticity loss in continual Atari setups.
@@
We use the CartPole-v1 and Acrobot-v1 environments from the Gym Classic Control suite, along with LunarLander-v2 from Box2d Control. To introduce distribution shifts, a method outlined by Mendez et al. (2020) involves periodically altering the environment dynamics. Although such distribution shifts pose only mild challenges for robust methods like PPO with Adam (Appendix C). We instead implement a more challenging form of distribution shift. Every 200 steps we perturb each observation dimension with random noise within a range of ±$\pm 2$, treating each perturbation phase as a distinct task.
Figure 6: Reward performance across CartPole, Acrobot, and LunarLander Gym Control tasks. Both Adam PPO and CReLU experience extreme plasticity loss, failing to recover after the initial distribution shift. Conversely, Trac PPO successfully avoids such plasticity loss, rapidly adapting when facing extreme distribution shifts.
Here (Figure 6), we notice a peculiar behavior after introducing the first distribution shift in both Adam PPO and CReLU: policy collapse. We describe this as an extreme form of loss of plasticity. Surprisingly, Trac PPO remains resistant to these extreme distribution shifts. As we see in the Acrobot experiment, Trac PPO shows minimal to no policy damage after the first few distribution shifts, whereas Adam PPO and CReLU are unable to recover a policy at all. We investigate if Trac’s behavior here indicates positive transfer in Appendix A. Across the three Gym Control environments, Trac PPO shows an average normalized improvement of 204.18% over Adam PPO and 1044.24% over CReLU (Table 1).